# Uncertainty Calibration for Counterfactual Propensity Estimation in Recommendation

### v2

@@ Abstract
Post-click conversion rate (CVR) is a reliable indicator of online customers’ preferences, making it crucial for developing recommender systems. A major challenge in predicting CVR is severe selection bias, arising from users’ inherent self-selection behavior and the system’s item selection process. To mitigate this issue, the inverse propensity score (IPS) is employed to weight the prediction error of each observed instance. However, current propensity score estimations are unreliable due to the lack of a quality measure. To address this, we evaluate the quality of propensity scores from the perspective of uncertainty calibration, proposing the use of expected calibration error (ECE) as a measure of propensity-score quality. We argue that the performance of IPS-based recommendations is hampered by miscalibration in propensity estimation. We introduce a model-agnostic calibration framework for propensity-based debiasing of CVR predictions. Theoretical analysis on bias and generalization bounds demonstrates the superiority of calibrated propensity estimates over uncalibrated ones. Experiments conducted on the Coat, Yahoo and KuaiRand datasets show improved uncertainty calibration, as evidenced by lower ECE values, leading to enhanced CVR prediction outcomes.
@@ I Introduction
Figure 1: For recommendation with MNAR on the Coat shopping dataset, we use the raw inverse propensity estimator with and without the platt scaling calibration and give the scatter plot of the expected propensity vs the fraction of observed ratings. The diagonal line is the perfect uncertainty calibration result. As can be seen, the raw propensity estimations are severely miscalibrated.
@@
In machine learning methods widely used in recommendation systems, uncertainty quantification is often poorly characterized, leading to over-confident predictions. This issue is prevalent not only in deep learning models [12] but also in shallow models, such as logistic regression [13]. The uncertainty of personalized ranking probabilities can be learned through uncertainty calibration methods [14, 15] and applied in online advertising systems [16, 17].
@@
Despite this, propensity scores are frequently miscalibrated, limiting the effectiveness of IPS, even though IPS has been validated in recommendation systems and other applications. As illustrated in Fig. 1, expected propensity scores are not calibrated with the fraction of observed samples, deviating from perfect calibration (the diagonal line). In terms of uncertainty calibration, expected propensity scores, such as 95%, should correspond to the same level of observed sample fraction (95%). The uncertainty originates from inaccurate propensity score predictions, leading to inaccurate recommendations when dealing with MNAR data using miscalibrated propensity scores.
In this paper, we propose using expected calibration error (ECE) as a quality measure for the reliability of propensity scores and highlight the miscalibration issue in current propensity estimation. Additionally, we introduce a model-agnostic uncertainty calibration framework for propensity estimation. Extensive experiments demonstrate that lower ECE values in propensity scores result in better CVR prediction outcomes.
The contribution of this paper are as follows:
  • We identify the miscalibration issue in propensity estimation and propose using ECE as a quality measure for the reliability of propensity scores.
  • We demonstrate that miscalibration of propensity scores limits the debiasing performance of propensity-based CVR prediction and propose a thoughtful uncertainty calibration methodology for propensity scores.
  • We provide theoretical analysis and experimental results to show the superiority of uncertainty-calibrated propensity scores for unbiased CVR prediction.
@@ II-A Propensity-based Debiasing Recommendation
Recently, the more robust doubly robust (MRDR) method [1] enhances the robustness of DR-JL by optimizing the variance of the DR estimator with the imputation model. Specifically, MRDR keeps the loss of the CVR prediction model in Eqn. 5 unchanged, which replacing the loss of the imputation model in Eqn. 6 with the following loss
@@ II-C Uncertainty Calibration for Deep Learning
Regarding calibration methodologies, one effective approach is Bayesian generative modeling, with representative models including Bayesian neural networks and deep Gaussian processes [33, 34]. Bayesian neural networks are generally computationally expensive to train, so approximate methods have been developed, such as MC-Dropout [23] and deep ensembles [35]. Alternatively, uncertainties can be obtained from the calibration of inaccurate uncertainties. Methods employing scaling and binning for calibration are used for both classification and regression models [36, 37, 12, 38, 39, 40, 41]. An additional advantage of calibration methods is their model-agnostic nature, making them applicable to any IPS-based model.
@@ III-B Uncertainty Calibration for Propensity Estimation
As illustrated in Fig. 1, raw propensity estimation models are generally miscalibrated, often producing overconfident probability predictions. To reduce biases and achieve calibrated propensity scores, we consider a model-agnostic uncertainty calibration $q$ in conjunction with the propensity learning model $g_{\phi}$. Specifically, two model-agnostic methods are considered for propensity probability calibration: 1) uncertainty probability quantification and 2) post-processing uncertainty calibration.
@@ III-B1 Uncertainty Probability Quantification for Propensity scores
Due to the challenges and high computational cost associated with fully Bayesian models, we propose two approximate uncertainty quantification methods for propensity estimation: 1) Monte Carlo Dropout [43] and 2) deep ensembles [35].
@@ III-B2 Post-processing Calibration for Propensity scores
where $g_{\phi}(x)$ represents the original propensity outputs, $\sigma$ is the sigmoid function, and $b,c$ are learnable parameters of the sigmoid function [36]. Platt scaling is equivalent to class-conditional Gaussian likelihoods with the same variance. For multi-class classification, Platt scaling can be augmented with a temperature parameter to soften the softmax output, known as temperature scaling [12]. In this paper, we employ Platt scaling for the NMAR binary setting.
@@ 17       Update $\theta$ by descending along the gradient $\nabla_{\theta}\mathcal{L}_{\textrm{IPS}}(\theta)$;
With the calibrated propensity scores, we can train the propensity-based CVR prediction debiasing model in two steps: first, train the propensity estimation model $g_{\phi}$, obtain the calibrated propensity scores, and then train the CVR prediction model $f_{\theta}$ using these inverse calibrated propensities. This process is detailed in Algorithm 1.
@@ III-B3 Computational Complexity Analysis
  • Post-Processing Calibration (e.g., Platt Scaling) involves adjusting the outputs of an already trained model using additional parameters (like $b$ and $c$ in Platt scaling). The primary computational expense here is the forward pass to compute $g_{\phi}(x)$ and the subsequent optimization to learn the calibration parameters. This can generally be much less computationally intensive compared to the previous methods, as it typically involves simpler operations over the model’s outputs and potentially fewer parameters to optimize.
@@ III-C Theoretical Analysis of Uncertainty Calibration using Expected Calibration Errors
By calibrating the propensity uncertainty, the expected calibration error can be reduced, leading to improved CVR predictions. We now provide a theoretical analysis of the proposed method.
@@
Theorem 2 reveals the bias-variance tradeoff in the real-world performance of the calibrated inverse propensity score estimator. A smaller bias results from reduced propensity bias. Additionally, lower propensity estimation error and expected calibration error (ECE) lead to better recommendation results.
We assume that the calibrated IPS has lower estimation error and calibration loss. Furthermore, IPS predictions are generally overestimated. With these two assumptions, we derive the following corollary.
@@
Compared with the inverse propensity score estimator, the prediction error bound of the calibrated doubly robust estimator has a smaller bias and has a upper bound that is propotional to ECE:
@@
Corollary 2 demonstrates that the expected calibration error (ECE) effectively bounds the final prediction error.
@@ IV-A Experimental Setting
  • Coat Shopping [^1] https://www.cs.cornell.edu/~schnabts/mnar/: The coat shopping dataset was collected to simulate the missing not at random data of user shopping for coats online. The costumers were ordered to find their favorite coats in the online store. After browsing, the users were asked to rate 24 coats they had explored before and 16 randomly picked ones on a five point scale. It contains ratings from 290 users of 300 items. There are 6960 MNAR ratings and 4640 MAR ratings.
@@
  • KuaiRand [^3] https://kuairand.com/: this datasets includes 7583 videos and 27285 users, containing 1436609 biased data and 1186509 unbiased data. Following recent work[46], we regard the biased dsata as the training set, and utilize the unbiased data for model validation (10%) and evaluation (90%).
@@
For uncertainty calibration, we assess the expected calibration error (ECE) using Eqn. 9.
For evaluating recommendation results, we employ two metrics: discount cumulative gain (DCG) and Recall. These are defined as:
${\rm DCG}(K)=\sum_{k=1}^{K}\frac{Rel_{k}}{\log_{2}(k+1)},$ (24)
${\rm Recall}(K)=\sum_{k=1}^{K}Rel_{k},$ (25)
Further implementation details, including evaluation metrics, optimization, and hyperparameters for all baselines, can be found in Section II in Supplemental Materials.
@@ IV-B Calibration Results of propensity scores(RQ1)
As shown in Table I, the calibration methods, especially Platt scaling, significantly reduce the Expected Calibration Error (ECE). Compared to uncalibrated propensity scores, Platt scaling reduces the ECE by more than a factor of three. Figure 2 presents the calibration curves and propensity histograms of the calibrated propensity scores, where ”Raw” denotes uncalibrated propensity scores. In Figure 2(a), calibration narrows the gap between the raw propensity model and the perfect propensity model (represented by the diagonal line). Since propensity-based debiasing methods rely solely on propensity scores from click events, the right side of the calibration curve further validates the effectiveness of propensity score calibration.
@@
(a) Calibration Curve
(b) propensity scores Histogram
Datasets Methods DCG@K Recall@K Average
@@
Coat Shopping ${\rm Neumf}_{base}$ 0.7478 1.0152 1.1989 0.8705 1.4435 1.9371 1.2021
Raw 0.7472$\pm$0.0143 1.0204$\pm$0.0124 1.2010$\pm$0.0111 0.8738$\pm$0.0176 1.4591$\pm$0.0177 1.9451$\pm$0.0227 1.2077
MC Dropout 0.7651$\pm$0.0242 1.0322$\pm$0.0172 1.2101$\pm$0.0145 0.8962$\pm$0.0283 1.4675$\pm$0.0266 1.9443$\pm$0.0209 1.2192
Deep Ensembles 0.7584$\pm$0.0271 1.0256$\pm$0.0232 1.2065$\pm$0.0225 0.8848$\pm$0.0342 1.4574$\pm$0.0341 1.9454$\pm$0.0322 1.2127
Platt Scaling 0.7627$\pm$0.0155 1.0405$\pm$0.0215 1.2259$\pm$0.0171 0.8890$\pm$0.0168 1.4823$\pm$0.0374 1.9806$\pm$0.0229 1.2301
Yahoo! R3 ${\rm Neumf}_{base}$ 0.5277$\pm$0.0209 0.7352$\pm$0.0209 0.8630$\pm$0.0209 0.6333$\pm$0.0209 1.0769$\pm$0.0209 1.4202$\pm$0.0209 0.8760
Raw 0.5433$\pm$0.0056 0.7395$\pm$0.0065 0.8669$\pm$0.0062 0.6468$\pm$0.0048 1.0661$\pm$0.0063 1.4086$\pm$0.0057 0.8785
MC Dropout 0.5410$\pm$0.0178 0.7406$\pm$0.0202 0.8663$\pm$0.0187 0.6452$\pm$0.0197 1.0720$\pm$0.0248 1.4094$\pm$0.0211 0.8790
Deep Ensembles 0.5342$\pm$0.0043 0.7412$\pm$0.0047 0.8692$\pm$0.0027 0.6404$\pm$0.0049 1.0831$\pm$0.0068 1.4270$\pm$0.0047 0.8825
Platt Scaling 0.5470$\pm$0.0065 0.7535$\pm$0.0033 0.8778$\pm$0.0039 0.6528$\pm$0.0088 1.0941$\pm$0.0053 1.4275$\pm$0.0060 0.8921
TABLE II: Overall IPS-based recommendation performance on Coat Shopping and Yahoo! R3. The best results are shown in boldface and the second best results are marked using underline.
Figure 2(b) shows the propensity histograms of the calibrated IPS methods trained on the Coat Shopping dataset. It can be observed that the calibrated propensity scores not only exhibit lower ECE but also demonstrate reduced polarization. Both ECE and polarization are crucial aspects of propensity scores. The calibration curve and propensity histograms for the Yahoo! R3 dataset are detailed in Figure 3, which supports the findings from Figure 2.

### v3

@@ Abstract
Post-click conversion rate (CVR) is a reliable indicator of online customers’ preferences, making it crucial for developing recommender systems. A major challenge in predicting CVR is severe selection bias, arising from users’ inherent self-selection behavior and the system’s item selection process. To mitigate this issue, the inverse propensity score (IPS) is employed to weight the prediction error of each observed instance. However, current propensity score estimations are unreliable due to the lack of a quality measure. To address this, we evaluate the quality of propensity scores from the perspective of uncertainty calibration, proposing the use of Expected Calibration Error (ECE) as a measure of propensity-score quality, which quantifies the extent to which predicted probabilities are overconfident by assessing the difference between predicted probabilities and actual observed frequencies. Miscalibrated propensity scores can lead to distorted IPS weights, thereby compromising the debiasing process in CVR prediction. In this paper, we introduce a model-agnostic calibration framework for propensity-based debiasing of CVR predictions. Theoretical analysis on bias and generalization bounds demonstrates the superiority of calibrated propensity estimates over uncalibrated ones. Experiments conducted on the Coat, Yahoo and KuaiRand datasets show improved uncertainty calibration, as evidenced by lower ECE values, leading to enhanced CVR prediction outcomes.
@@ I Introduction
Figure 1: For recommendation with MNAR on the Coat shopping dataset, we use the raw propensity estimator with and without the platt scaling calibration and give the scatter plot of the expected propensity vs the fraction of observed ratings. The diagonal line is the perfect uncertainty calibration result. As can be seen, the raw propensity estimations are severely miscalibrated.
@@
In machine learning methods widely used in recommendation systems, uncertainty quantification is often poorly characterized, leading to over-confident predictions. This issue is prevalent not only in deep learning models [12] but also in shallow models, such as logistic regression [13]. Both types of models are prone to overconfidence because they are typically optimized to minimize error metrics without explicitly accounting for uncertainty, resulting in predictions that underestimate the true uncertainty in the data [12]. The uncertainty of personalized ranking probabilities can be learned through uncertainty calibration methods [14, 15] and applied in online advertising systems [16, 17].
@@
Calibration in machine learning refers to the degree to which predicted probabilities reflect the true likelihood of an event. A model is considered well-calibrated if, for predictions assigned a probability of $p$, the actual frequency of the event is also $p$. The propensity scores are frequently miscalibrated, limiting the effectiveness of IPS, even though IPS has been validated in recommendation systems and other applications. As illustrated in Fig. 1, expected propensity scores are not calibrated with the fraction of observed samples, deviating from perfect calibration (the diagonal line). Please note that the results in Fig. 1 also applies to the more recent methods, DR-JL [10] and MRDR [1]. In terms of uncertainty calibration, expected propensity scores, such as 95%, should correspond to the same level of observed sample fraction (95%). The uncertainty originates from inaccurate propensity score predictions, leading to inaccurate recommendations when dealing with MNAR data using miscalibrated propensity scores.
In machine learning methods commonly used in recommendation systems, both deep and shallow models are prone to overconfidence, often underestimating the true uncertainty in the data. This overconfidence stems from optimizing error metrics without explicitly accounting for uncertainty, which can lead to miscalibrated predictions [12, 13]. Propensity scores, crucial for IPS-based debiasing, are particularly susceptible to this issue, as demonstrated in Fig. 1, where the observed miscalibration deviates significantly from perfect uncertainty calibration. Miscalibrated propensity scores result in distorted IPS weights, ultimately compromising the debiasing process and prediction reliability. To address this, we propose using Expected Calibration Error (ECE) as a robust measure of propensity-score quality. By quantifying the degree of miscalibration, ECE reveals the extent to which overconfidence in propensity predictions hampers IPS effectiveness. Lower ECE values indicate better-calibrated scores, enabling more accurate and unbiased CVR predictions. This highlights the critical importance of addressing miscalibration to enhance the reliability and robustness of IPS-based methods in handling MNAR data.
The contributions of this paper are as follows:
  • Identification of Propensity Miscalibration: We reveal a critical issue in current CVR prediction approaches—the miscalibration of propensity scores—and propose Expected Calibration Error (ECE) as a robust metric to assess the reliability of these scores.
  • Uncertainty Calibration Framework: We introduce a novel, model-agnostic framework for uncertainty calibration of propensity scores, significantly improving their reliability and aligning them more closely with observed data distributions.
  • Enhanced Debiasing Effectiveness: By addressing miscalibration, we demonstrate how calibrated propensity scores bolster the debiasing performance of inverse propensity score (IPS)-based and doubly robust (DR) learning methods in recommendation systems.
  • Comprehensive Theoretical and Empirical Validation: We provide rigorous theoretical analysis of bias reduction and generalization bounds, supported by extensive experiments on benchmark datasets (e.g., Coat, Yahoo, and KuaiRand). Results highlight the superiority of calibrated propensity scores in achieving accurate and unbiased CVR predictions, as evidenced by lower ECE values and improved prediction performance.
@@ II-A Propensity-based Debiasing Recommendation
Recently, the more robust doubly robust (MRDR) method [1] enhances the robustness of DR-JL by optimizing the variance of the DR estimator with the imputation model. Specifically, MRDR keeps the loss of the CVR prediction model in Eqn. 5 unchanged, which replaces the loss of the imputation model in Eqn. 6 with the following loss
@@ II-C Uncertainty Calibration for Deep Learning
Regarding calibration methodologies, one effective approach is Bayesian generative modeling, with representative models including Bayesian neural networks and deep Gaussian processes [33, 34]. Bayesian neural networks are generally computationally expensive to train, so approximate methods have been developed, such as MC-Dropout [23] and deep ensembles [35]. Alternatively, uncertainties can be obtained from the calibration of inaccurate uncertainties. Methods employing scaling and binning for calibration are used for both classification and regression models [36, 37, 12, 38, 39, 40, 41, 42, 43, 44]. An additional advantage of calibration methods is their model-agnostic nature, making them applicable to any IPS-based model. However, the joint modeling of the model calibration and the base model, namely the CVR model in this paper, might bring synergy between the two tasks. The possible way includes the fully Bayesian generative modeling of the two parts or transfrom the uncertainty calibration as a loss regularization term in the CVR modeling, where both faces several challenges for modeling and training.
@@ III-B Uncertainty Calibration for Propensity Estimation
As illustrated in Fig. 1, raw propensity estimation models are generally miscalibrated, often producing overconfident probability predictions. The reason of the overconfident and miscalibrated propensity estimation models is that they are typically optimized to minimize error metrics without explicitly accounting for uncertainty, which shares the same reason of the overconfidence of the conventional deep models. To reduce biases and achieve calibrated propensity scores, we consider a model-agnostic uncertainty calibration $q$ in conjunction with the propensity learning model $g_{\phi}$. Specifically, two model-agnostic methods are considered for propensity probability calibration: 1) uncertainty probability quantification and 2) post-processing uncertainty calibration.
@@ III-B1 Uncertainty Probability Quantification for Propensity scores
Due to the challenges of performing exacting inference and its high computational cost associated with fully Bayesian models, we propose two approximate uncertainty quantification methods for propensity estimation: 1) Monte Carlo Dropout [46] , 2) deep ensembles [35] and 3) dual focal loss [47].
@@
Dual Focal Loss[47] not only considers the ground truth logit, but also take into account the highest logit ranked after the ground truth logit. By maximizing the gap between these two logits, our dual focal loss can achieve a better balance between over-confidence and under-confidence.
@@ III-B2 Post-processing Calibration for Propensity scores
where $g_{\phi}(x)$ represents the original propensity outputs, $\sigma$ is the sigmoid function, and $b,c$ are learnable parameters of the sigmoid function [36]. With the goal of achieving better alignment between predicted and true probabilities, parameters $b$ and $c$ are optimized using a negative log-likelihood (NLL) objective on a held-out calibration dataset. This process enables the model to learn a calibration mapping that minimizes overconfidence or underconfidence in predictions, thereby enhancing overall calibration performance. Platt scaling is equivalent to class-conditional Gaussian likelihoods with the same variance. For multi-class classification, Platt scaling can be augmented with a temperature parameter to soften the softmax output, known as temperature scaling [12]. In this paper, we employ Platt scaling for the NMAR binary setting.
@@ 17       Update $\theta$ by descending along the gradient $\nabla_{\theta}\mathcal{L}_{\textrm{IPS}}(\theta)$;
With the calibrated propensity scores, we can train the propensity-based CVR prediction debiasing model in two steps: first, train the propensity estimation model $g_{\phi}$, obtain the calibrated propensity scores, and then train the CVR prediction model $f_{\theta}$ using these inverse calibrated propensities. This process is detailed in Algorithm 1. Please note that, if the calibration method is Dual Focal loss, the loss gradient in line 4 would consider the added dual focal loss and no other calibration step (line 5-8) is needed.
Our proposed method is built upon but differs from existing calibration approaches in several key ways. First, it is model-agnostic, enabling application across various propensity-based models without altering their underlying architectures. Second, unlike traditional calibration methods focused on general classification tasks, our approach specifically addresses propensity score calibration for post-click conversion rate (CVR) prediction, tackling the selection bias inherent in recommendation systems. Third, it uniquely integrates with debiasing techniques such as Inverse Propensity Scoring (IPS) and Doubly Robust (DR) learning, enhancing their effectiveness by resolving miscalibrated propensity scores.
Our method is built upon the exsting calibration approachs but stands out by being model-agnostic, specifically targeting propensity score calibration for CVR prediction to address selection bias in recommendation systems. Unlike existing methods, it integrates with debiasing techniques like IPS and DR learning, improving their effectiveness. Additionally, it prioritizes Expected Calibration Error (ECE) as a key metric, offering a focused evaluation of calibration quality in the context of the counterfactual propensity estimation in recommendation.
@@ III-B3 Computational Complexity Analysis
  • Post-Processing Calibration (e.g., Platt Scaling) involves adjusting the outputs of an already trained model using additional parameters (like $b$ and $c$ in Platt scaling). The primary computational expense here is the forward pass to compute $g_{\phi}(x)$ and the subsequent optimization to learn the calibration parameters. This can generally be much less computationally intensive compared to the previous methods, as it typically involves simpler operations over the model’s outputs and potentially fewer parameters to optimize. Dual Focal Loss optimizes both the ground truth logit and the highest competing logit by maximizing the gap between them. Since it modifies the loss function during training, the computational complexity is comparable to standard training with additional gradient computations to handle the second logit. It remains computationally less expensive than MC Dropout or Deep Ensembles since it does not require multiple forward passes or model ensembles.
@@ III-C Theoretical Analysis of Uncertainty Calibration using Expected Calibration Errors
The miscalibration of propensity scores, driven by overconfidence in both deep and shallow models, distorts IPS weights and hampers CVR predictions. By using Expected Calibration Error (ECE) to measure miscalibration, we demonstrate that reducing ECE improves propensity reliability and enhances prediction accuracy. By calibrating the propensity uncertainty, the Expected Calibration Error can be reduced, leading to improved CVR predictions. We now provide a theoretical analysis of the proposed method.
@@
Theorem 2 reveals the bias-variance tradeoff in the real-world performance of the calibrated inverse propensity score estimator. A smaller bias results from reduced propensity bias.
Based theorem 2, better recommendation results come from only lower propensity estimation error but also Expected Calibration Error (ECE). Therein, lower ECE is the main goal of our uncertainty calibration algorithm since IPS predictions are generally overestimated. Therefore, with these two assumptions, we derive the following corollary.
@@
Compared with the inverse propensity score estimator, the prediction error bound of the calibrated doubly robust estimator has a smaller bias and has a upper bound that is proportional to ECE:
@@
Corollary 2 demonstrates that the Expected Calibration Error (ECE) effectively bounds the final prediction error. Several existing works give proper theoretical anlaysis for the uncertainty calibration works, including parametric calibration function [13] and unparametric binning method [50]. These works reinforce the soundness of our approach and provide a broader theoretical context for uncertainty calibration.
@@ IV-A Experimental Setting
  • Coat Shopping [^1] https://www.cs.cornell.edu/~schnabts/mnar/: The coat shopping dataset was collected to simulate the missing not at random data of user shopping for coats online. The customers were ordered to find their favorite coats in the online store. After browsing, the users were asked to rate 24 coats they had explored before and 16 randomly picked ones on a five point scale. It contains ratings from 290 users of 300 items. There are 6960 MNAR ratings and 4640 MAR ratings.
@@
  • KuaiRand [^3] https://kuairand.com/: this datasets includes 7583 videos and 27285 users, containing 1436609 biased data and 1186509 unbiased data. Following recent work[51], we regard the biased data as the training set, and utilize the unbiased data for model validation (10%) and evaluation (90%).
@@
  • Dual Focal Loss[47]: We implement the Dual Focal Loss with a gamma parameter of 2.0 to train the propensity model. The loss function considers both the probability of the target label and the highest probability among all other labels that are smaller than the target probability, which helps to achieve better calibration.
@@
For uncertainty calibration, we assess the Expected Calibration Error (ECE) using Eqn. 9. Other evaluation metrics include AUC, discount cumulative gain (DCG) and Recall [1]. Further implementation details, including evaluation metrics, optimization, and hyperparameters for all baselines, can be found in Section II in Supplemental Materials.
@@ IV-B Calibration Results of propensity scores(RQ1)
As shown in Table I, the calibration methods, especially Platt scaling, significantly reduce the Expected Calibration Error (ECE). Compared to uncalibrated propensity scores, Platt scaling reduces the ECE by more than a factor of three. Figure 2 presents the calibration curves and propensity histograms of the calibrated propensity scores, where “Raw” denotes uncalibrated propensity scores. In Figure 2(a), calibration narrows the gap between the raw propensity model and the perfect propensity model (represented by the diagonal line). Since propensity-based debiasing methods rely solely on propensity scores from click events, the right side of the calibration curve further validates the effectiveness of propensity score calibration.
@@
Datasets Methods AUC DCG@K Recall@K
@@
Coat Shopping ${\rm Neumf}_{base}$ 0.7604$\pm$0.0041 0.7478$\pm$0.0201 1.0152$\pm$0.0222 1.1989$\pm$0.0243 0.8705$\pm$0.0287 1.4435$\pm$0.0301 1.9371$\pm$0.0205
Raw 0.7578$\pm$0.0036 0.7472$\pm$0.0143 1.0204$\pm$0.0124 1.2010$\pm$0.0111 0.8738$\pm$0.0176 1.4591$\pm$0.0177 1.9451$\pm$0.0227
MC Dropout 0.7632$\pm$0.0011 0.7651$\pm$0.0242 1.0322$\pm$0.0172 1.2101$\pm$0.0145 0.8962$\pm$0.0283 1.4675$\pm$0.0266 1.9443$\pm$0.0209
Deep Ensembles 0.7675$\pm$0.0059 0.7584$\pm$0.0271 1.0256$\pm$0.0232 1.2065$\pm$0.0225 0.8848$\pm$0.0342 1.4574$\pm$0.0341 1.9454$\pm$0.0322
Dual FocalLoss 0.7611$\pm$0.0029 0.7491$\pm$0.0154 1.0264$\pm$0.0129 1.2034$\pm$0.0132 0.8751$\pm$0.0253 1.4458$\pm$0.0214 1.9479$\pm$0.0299
Platt Scaling 0.7693$\pm$0.0036 0.7627$\pm$0.0155 1.0405$\pm$0.0215 1.2259$\pm$0.0171 0.8890$\pm$0.0168 1.4823$\pm$0.0374 1.9806$\pm$0.0229
Yahoo! R3 ${\rm Neumf}_{base}$ 0.7131$\pm$0.0009 0.5277$\pm$0.0209 0.7352$\pm$0.0209 0.8630$\pm$0.0209 0.6333$\pm$0.0209 1.0769$\pm$0.0209 1.4202$\pm$0.0209
Raw 0.7172$\pm$0.0034 0.5433$\pm$0.0056 0.7395$\pm$0.0065 0.8669$\pm$0.0062 0.6468$\pm$0.0048 1.0661$\pm$0.0063 1.4086$\pm$0.0057
MC Dropout 0.7216$\pm$0.0027 0.5410$\pm$0.0178 0.7406$\pm$0.0202 0.8663$\pm$0.0187 0.6452$\pm$0.0197 1.0720$\pm$0.0248 1.4094$\pm$0.0211
Deep Ensembles 0.7254$\pm$0.0010 0.5342$\pm$0.0043 0.7412$\pm$0.0047 0.8692$\pm$0.0027 0.6404$\pm$0.0049 1.0831$\pm$0.0068 1.4270$\pm$0.0047
Dual FocalLoss 0.7219$\pm$0.0058 0.5441$\pm$0.0114 0.7452$\pm$0.0131 0.8720$\pm$0.0120 0.6515$\pm$0.0119 1.0708$\pm$0.0147 1.4216$\pm$0.0115
Platt Scaling 0.7235$\pm$0.0017 0.5470$\pm$0.0065 0.7535$\pm$0.0033 0.8778$\pm$0.0039 0.6528$\pm$0.0088 1.0941$\pm$0.0053 1.4275$\pm$0.0060
TABLE II: Overall IPS-based recommendation performance on Coat Shopping and Yahoo! R3. The best results are shown in boldface, and the second best results are marked using underline. ‡ indicates statistically significant improvements over the Raw method at p $<$ 0.05 level.
Figure 2(b) shows the propensity histograms of the calibrated IPS methods trained on the Coat Shopping dataset. It can be observed that the calibrated propensity scores not only exhibit lower ECE but also demonstrate reduced polarization. Both ECE and polarization are crucial aspects of propensity scores. The calibration curve and propensity histograms for the Yahoo! R3 dataset are detailed in Figure 2 in the Section IV in the Supplemental Material, which supports the findings from Figure 2.